OpenStudy (anonymous):

(i) Factorise then solve the equation x2 + 2x − 24 = 0. Help????

OpenStudy (anonymous):

Again step by step would be spiffing lol

OpenStudy (anonymous):

2x + 2x - 24 =0 4x -24 =0 +24 +24 4x = 24 Divide by 4. x= 6

OpenStudy (anonymous):

oh i'm such a twit can't believe i didn't get that it's cause it wants me to factorise it which confused me

OpenStudy (mathmale):

That "factorise" is a fancy word which I'd replace by "factor" in explaining this problem. With that " s " in the spelling of "factorise," I'm wondering whether this is British English.